# Axioms for AI Alignment from Human Feedback

### v1

@@ 7 $c^{-}_{1}$ $c^{-}_{2}$ $c^{-}_{3}$
We argue the ranking $a^{*}\succ c^{+}_{1}\succ c^{+}_{2}\succ c^{+}_{3}\succ c^{-}_{3}\succ c^{-}_{2}\succ c^{-}_{1}$ is a PMC ranking and but no linear reward function can position $a^{*}$ at the top of this ranking.
@@
Given the placement of $c^{\pm}_{i}$ at the standard basis vectors, each $c_{i}^{\pm}$ achieves a reward equivalent to one of the absolute values of the components of $\theta$, thus surpassing $r_{a^{*}}^{\theta}$ since:
$r_{a^{*}}^{\theta}<\max_{i}|\theta_{i}|$
Appendix B Additional Axiomatic Properties of Social Choice-Based Rules
@@
B.1 Copeland Violates Separability
Theorem B.2.
@@ B.1 Copeland Violates Separability
$v_{1}:a\succ g\succ d\succ e\succ f\succ b\succ c$
$v_{2}:b\succ a\succ c\succ e\succ d\succ g\succ f$
$v_{3}:b\succ a\succ e\succ d\succ f\succ g\succ c$
$v_{4}:c\succ e\succ f\succ g\succ b\succ a\succ d$
$v_{5}:d\succ c\succ f\succ g\succ b\succ e\succ a$
and
$v_{6}:a\succ d\succ b\succ f\succ c\succ g\succ e$
$v_{7}:b\succ a\succ e\succ d\succ c\succ g\succ f$
$v_{8}:c\succ a\succ e\succ b\succ f\succ g\succ d$
One can check that in the first profile, the Copeland scores in the first profile are $5,4,3,3,3,2,1$ for candidates $a,b,c,d,e,f,g$, respectively. Similarly, in the second profile, they are $6,5,3,3,3,1,0$. So under any consistent tie-breaking rule, both of these profiles would output the ranking $a\succ b\succ c\succ d\succ e\succ f\succ g$.
@@
B.2 Linear Kemeny Rule
@@
Note that Kemeny is in general NP-hard to compute. However, even for linear Kemeny, there is at least an exponential time aglorithm by brute-force computing the score of every ranking, and determining whether or not it is feasible.
Theorem B.3.
@@ 20 $(2,0,0,0,0,0,0)$
Each vector is constructed such that candidates are prioritized based on the magnitude of their first non-zero entry, leading to a natural ordering within grouped subsets: $\{1,2\}\succ\{3,4,5\}\succ\{6,7,8\}\succ\{9,10,11\}\succ\{12,13,14\}\succ\{15,16,17\}\succ\{18,19,20\}$. We will have six voters, with rankings induced by the parameter vectors described in Table 4.
Table 4: Paramater vectors of voters
Voter Preference Vector
@@
First, consider a set of candidates and a profile similar to the one that is given in the proof of Theorem B.4. When we restrict to reward functions that $1$ above $2$, then we can check that Linear Kemeny subject to PO outputs one of the rankings that are in the input profile. Without loss of generality, assume that it outputs the ranking of $v_{1}$.
Second, consider the same set of candidates and three voters, with rankings induced by the parameter vectors described in Table 5.
Table 5: Paramater vectors of voters
Voter Preference Vector
@@
B.3 Leximax Plurality
Plurality is probably the most ubiquitous voting rule in the world. Its ranking variant ranks the candidates in decreasing order with respect to their plurality scores. The plurality score of a candidate is equal to the number of her appearances in the first position. This rule is known to satisfy several axioms but in linear social choice cannot be directly applied, since not all rankings are feasibly.
Similarly to leximax Copeland, we define leximax Plurality as follows. It ranks first the candidate with the highest plurality score that can be ranked first under some parameter vector. Subject to this first position, it ranks second the candidate with the highest plurality score that can be ranked second under some parameter vector, and so on until all the positions are filled.
Theorem B.6.
@@ B.3 Leximax Plurality
Note that leximax Plurality always returns a ranking in which the candidate with the highest plurality score is ranked first, since there exists at least one feasibly ranking in which this candidate is ranked first. From this observation, we immediately see that leximax Plurality satisfies majority consistency.

### v2

@@ 7 $c^{-}_{1}$ $c^{-}_{2}$ $c^{-}_{3}$
We argue the ranking $a^{*}\succ c^{+}_{1}\succ c^{+}_{2}\succ c^{+}_{3}\succ c^{-}_{3}\succ c^{-}_{2}\succ c^{-}_{1}$ is a PMC ranking but no linear reward function can position $a^{*}$ at the top of this ranking.
@@
Given the placement of $c^{\pm}_{i}$ at the standard basis vectors, each $c_{i}^{\pm}$ achieves a reward equivalent to one of the absolute values of the components of $\theta$, thus surpassing $r_{a^{*}}^{\theta}$ since
$r_{a^{*}}^{\theta}<\max_{i}|\theta_{i}|.$
Appendix C Additional Axiomatic Properties of Social Choice-Based Rules
@@
C.1 Copeland Violates Separability
Theorem C.2.
@@ C.1 Copeland Violates Separability
Rank $v_{1}$ $v_{2}$ $v_{3}$ $v_{4}$ $v_{5}$
1 $a$ $b$ $b$ $c$ $d$
2 $g$ $a$ $a$ $e$ $c$
3 $d$ $c$ $e$ $f$ $f$
4 $e$ $e$ $d$ $g$ $g$
5 $f$ $d$ $f$ $b$ $b$
6 $b$ $g$ $g$ $a$ $e$
7 $c$ $f$ $c$ $d$ $a$
Rank $v_{6}$ $v_{7}$ $v_{8}$
1 $a$ $b$ $c$
2 $d$ $a$ $a$
3 $b$ $e$ $e$
4 $f$ $d$ $b$
5 $c$ $c$ $f$
6 $g$ $g$ $g$
7 $e$ $f$ $d$
In the first profile, the Copeland scores are $5,4,3,3,3,2,1$ for candidates $a,b,c,d,e,f,g$, respectively. Similarly, in the second profile, they are $6,5,3,3,3,1,0$. So under any consistent tie-breaking rule, both of these profiles would output the ranking $a\succ b\succ c\succ d\succ e\succ f\succ g$.
@@ 7 $e$ $f$ $d$
C.2 Linear Kemeny Rule
@@
Note that Kemeny is generally NP-hard to compute; however, even for linear Kemeny, there is at least an exponential time aglorithm by brute-force computing the score of every ranking, and determining whether or not it is feasible.
Theorem C.3.
@@ 20 $(2,0,0,0,0,0,0)$
Each vector is constructed such that candidates are prioritized based on the magnitude of their first non-zero entry, leading to a natural ordering within grouped subsets: $\{1,2\}\succ\{3,4,5\}\succ\{6,7,8\}\succ\{9,10,11\}\succ\{12,13,14\}\succ\{15,16,17\}\succ\{18,19,20\}$. We will have six voters, with rankings induced by the parameter vectors described below.
Voter Parameter Vector
@@
First, consider a set of candidates and a profile similar to the one that is given in the proof of Theorem C.4. When we restrict to reward functions that $1$ above $2$, then we can check that Linear Kemeny subject to PO outputs one of the rankings that are in the input profile. Without loss of generality, assume that it outputs the ranking of $v_{1}$.
Second, consider the same set of candidates and three voters, with rankings induced by the following parameter vectors:
Voter Parameter Vector
@@
C.3 Leximax Plurality
Plurality is probably the most ubiquitous voting rule in the world. Its ranking variant ranks the candidates in decreasing order with respect to their plurality scores. The plurality score of a candidate is equal to the number of her appearances in the first position. This rule is known to satisfy several axioms but in linear social choice cannot be directly applied, as not all rankings are feasibly.
Similarly to leximax Copeland, we define Leximax Plurality as follows. It ranks first the candidate with the highest plurality score that can be ranked first under some parameter vector. Subject to this first position, it ranks second the candidate with the highest plurality score that can feasibly be ranked second, and so on, until all the positions are filled.
Theorem C.6.
@@ C.3 Leximax Plurality
Note that leximax Plurality always returns a ranking in which the candidate with the highest plurality score is ranked first, since there exists at least one feasible ranking in which this candidate is ranked first. From this observation, we immediately see that leximax Plurality satisfies majority consistency.